OpenStudy (naveenbhatia1312):

Find an equation for the nth term of the arithmetic sequence. a19 = -92, a20 = 6

OpenStudy (mightychondrion):

Just from these two numbers you can find your common difference, then work backwards to find term 1 and write the equation.

OpenStudy (naveenbhatia1312):

@MightyChondrion i am a little confused with what you said or how to go about this

OpenStudy (mightychondrion):

In arithmetic sequences, we have this thing called the common difference. It's a number that tells us what we add to a term to get the next term. In this case, your CD is the number added to -92 to get 6, what is that number?

OpenStudy (naveenbhatia1312):

98

OpenStudy (mightychondrion):

Yeah, that's right. Do you know the formula for finding the nth term of an arithmetic sequence?

OpenStudy (naveenbhatia1312):

no

OpenStudy (naveenbhatia1312):

These are the four choices I was given an = -1856 + 98(n - 1) an = -1856 - 98(n - 1) an = -1856 - 98(n + 1) an = -1856 + 98(n + 1)

OpenStudy (mightychondrion):

\[u _{n}=u _{1}+(n-1)d\] Where \[u _{n}\] is your nth term, in this case, you have term 19, and u\[_{1}\] is the first term you need to find. N is simply what number term you have (19) and d is your common difference aka the number added to a term to get the next term.

OpenStudy (mightychondrion):

Now if you plug into this formula, you have \[u _{n}=-92\] so you can write that as \[-92=u _{1}+(19-1)*98\]. Solve for the first term now

OpenStudy (naveenbhatia1312):

-u1=1856

OpenStudy (naveenbhatia1312):

@MightyChondrium

OpenStudy (mightychondrion):

Mhm, now you can represent the nth term as \[u _{n}=-1856+98(n-1)\]